hi i got this question based on exponential methods i have tried but sadly not succeeding i keep getting answers that make no sense. the dacay of a radioactive substance is governed by the relationship N=N0e^ -kt
for a certain substance 20% has decayed in 23 years, what is the half life of this substance and how long will it take 89% of the substance to decay ?

Answer by scott8148(6628)

(1 - .2)No = No e^(-kt) ___ .8 = e^(-k * 23)

ln(.8) = -23k ___ ln(.8) / (-23) = k

solve for k (the decay constant), and then substitute to find t for 50% (half-life) and 89%